Cole is about to purchase 4 units of good A and 6 units of good B. The price of A is $3 and the price of B is $2. Cole has only $24 to spend. Assume that the marginal utility of the fourth unit of A is 18 and the marginal utility of the sixth unit of B is 6. If Cole wants to maximize utility he should buy less of B and more of A. Cole cannot maximize utility with only $24 he should buy less of A and more of B he should buy only B because it is cheaper he should buy A and B in the quantities indicated QUESTION 5 1 points Quantit Total Utility from X 24 42 56 Total Utility from Y 85 130 160 185 200 210 215 4 74 80 84 7 Refer to the table above. Assume that the price of good X is $2 per unit, the price of good Y is $5 per unit, and that the consumer spends a total of $44 on both goods. What combination of X and Y will the consumer purchase in order to maximize utility? A We dont have enough information of determine his consumption bundle He will spend more on X than on Y because X is less expensive thanY He will purchase 4 units of Y and 4 units of X. He will purchase 2 units of Y and 2 units of X. He will purchase 6 units of Y and 7 units of X
Answer #1

1) MU/P of A = 18/3 = 6

MU/P fo B = 6/2 =3

He should buy more of A and less of B

option(A)

2)

Q TU MU MU/P Q TU MU MU/P
1 24 24 12 1 85 85 17
2 42 18 9 2 130 45 9
3 56 14 7 3 160 30 6
4 66 10 5 4 185 25 5
5 74 8 4 5 200 15 3
6 80 6 3 6 210 10 2
7 84 4 2 7 215 5 1

Budget = 44

He will buy 7 units of X and 6 units of Y

option(E)